OpenStudy (anonymous):

Find the quotient x+7 / x / 9/ x

OpenStudy (ipwnbunnies):

\[\frac{x+7}{x} \div \frac{9}{x}\] Is that your problem?

OpenStudy (anonymous):

yes

OpenStudy (ipwnbunnies):

Do you know how to divide fractions, or a simple rule on how to do it?

OpenStudy (anonymous):

not really my teacher didnt really do good job on teacher us

OpenStudy (ipwnbunnies):

Kewl. B) Ok. Well. Remember this. Keep. Change. Flip. :3 Keep the first fraction, change the division sign to multiplication, flip the 2nd fraction.

OpenStudy (ipwnbunnies):

So, \[\frac{x+7}{x} \div \frac{9}{x}\] becomes \[\frac{x+7}{x} * \frac{x}{9}\]

OpenStudy (anonymous):

can you do cross multiplication

OpenStudy (anonymous):

or straight across

OpenStudy (ipwnbunnies):

No need to do cross multiplication here. These are two fractions not separated by an equal sign. Yes, do straight across multiplication. However, it's easier to look at it this way. You have an 'x' factor in both the numerator and denominator. You can already cancel those 'x's out. c: Now, what will you be left with?

OpenStudy (anonymous):

7x over 9

OpenStudy (anonymous):

or x+7/9

OpenStudy (ipwnbunnies):

o.o Nooo. How'd you get 7x? Yesh, (x+7). I'll show you the steps in btwn real quick.

OpenStudy (ipwnbunnies):

\[\frac{(x+7)(x)}{(x)(9)} = \frac{x+7}{9}\]

OpenStudy (anonymous):

thank you

OpenStudy (ipwnbunnies):

No prob c: